OpenStudy (alinaa1024):

Which function has the greatest rate of change on the interval from x = 3 pi over 2 to x = 2π? f(x) g(x) h(x) All three functions have the same rate of change.

OpenStudy (alinaa1024):

f(x) x y 0 0 pi/2 2 π 0 3 pi/2 −2 2π 0

OpenStudy (alinaa1024):

h(x) = 6 sin(x) + 1

OpenStudy (alinaa1024):

@ThereisnoUsername

OpenStudy (alinaa1024):

@Mehek14

OpenStudy (alinaa1024):

@agent0smith @satellite73

OpenStudy (alinaa1024):

@ThereisnoUsername

OpenStudy (peachpi):

\[rate~~of~~change = \frac{ f(b)-f(a) }{ b-a }\] for each of those a = 3π/2 and b = 2π Use the graph, formula, and chart to find the y-values for 3π/2 and 2π then put them in the formula

OpenStudy (alinaa1024):

i dont understand

OpenStudy (peachpi):

look at the chart you put up for f(x). What is y when x is 2π?

OpenStudy (alinaa1024):

0 @peachpi

OpenStudy (peachpi):

yes. and what if y when x is 3π/2

OpenStudy (alinaa1024):

-2 @peachpi

OpenStudy (mathmale):

Is this algebra for calculus? What does "rate of change" mean to you? What does "average rate of change" mean to you? To compare these 3 functions, you'll no doubt have to calculate the rate of change of each one on the specified interval.

OpenStudy (mathmale):

One thing I'd strongly suggest you do is to graph all three of the given functions on the specified interval. Then, by eye, determine which function has the greatest rate of change on that interval. You don't need calculus if you approach the problem in this way.